Two short-eared pigs are mated. Of the resulting the progeny three have no ears, seven have short ears, and four have long ears. Explain these results by diagramming the cross.
Answer #1

Given that, mating parents are both short eared, and they produce offspring which comprise 3 without ears, 7 with short ears and 4 with long ears.

In the progeny, ratio of no ear to short ear to long ears can be found by dividing each by 3 (value for phenotype with least number ):

3/3 = 1

7/3= 2.33

4/3= 1.33

It is approximately 1:2:1

This ratio corresponds to incomplete dominance, in which none of the allele is dominant when they are present together. This can be explained as follows:

Let the allele for no ear be "a" and allele for long ears be "A". An individual has two alleles for a gene, and for ear length we can have three possibilities: aa individuals, Aa individuals and AA individuals.

aa types will have no ears, AA types will have long ears.

But, Aa individuals will have a blend of the two alleles, A and a , and phenotype will be short ears.

Following the question, cross can be shown as:

Short ear (Aa) x Short ear (Aa) ( parents)

A a
A AA (long ear) Aa (short ear )
a Aa (short ear) aa (no ear)

Thus, we can see that Aa, Aa and aa will be in the ratio 1:2:1.